OpenStudy (anonymous):

If \(x^2 + y^2 = 1^2\) \(y' = -\)\(/\) (for \(y \neq 0\) ) \(y'' = -\)\(/y^3\)

OpenStudy (anonymous):

are you getting the question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i have to find single and double derivative of the given function

OpenStudy (anonymous):

first derivative :2x+2y=0 second, 2+2=4

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

i know how to find the first derivative tell me about the second

OpenStudy (anonymous):

ok 2x + 2y * dy/dx = 0 dy/dx = -2x / 2y = -x/y

OpenStudy (anonymous):

what about the second derivative?

myininaya (myininaya):

use the quotient rule

OpenStudy (anonymous):

i tried but no success

myininaya (myininaya):

\[y''=\frac{(-x)'y-(-x)y'}{y^2}\]

myininaya (myininaya):

(-x)'=-1

OpenStudy (anonymous):

second derivative use quotient rule y * -1 - (-x) * dy/dx ------------------- y^2 -y + x*dy/dx ---------- y^2

myininaya (myininaya):

replace y' with -x/y

myininaya (myininaya):

multiply top and bottom by y

myininaya (myininaya):

you will be able to use x^2+y^2=1 to simplify further

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

thanks myin - i got a bit lost after the second line!

OpenStudy (anonymous):

guys the answer im getting is wrong! :( can you solve it for me?

myininaya (myininaya):

\[y''=\frac{(-x)'y-(-x)y'}{y^2}=\frac{-1y+xy'}{y^2}=\frac{-y+x \cdot \frac{-x}{y}}{y^2}\] \[=\frac{-y \cdot y+x(-x)}{y \cdot y^2}=\frac{-y^2-x^2}{y^3}=\frac{-(x^2+y^2)}{y^3}\]

myininaya (myininaya):

now remember x^2+y^2=1

myininaya (myininaya):

\[=\frac{-1}{y^3}\]

OpenStudy (anonymous):

very nice....

OpenStudy (anonymous):

nice